OpenStudy (anonymous):

A particular sound wave can be graphed using the function y = 1 sin 2x. Find the amplitude of the function. A. amplitude = 1 B. amplitude = p C. amplitude = p D. amplitude = –1

OpenStudy (anonymous):

For any sinusoidal function given in the form:\[\bf y=Asin(bx-c)+K\]The amplitude is:\[\bf \ Amplitude=\left| A \right|\]

OpenStudy (anonymous):

I don't get it?

OpenStudy (anonymous):

It would be A.

OpenStudy (anonymous):

k thanks

OpenStudy (radar):

Just remember that the sin x or sin 2x can only range from -1 to 1.

OpenStudy (anonymous):

oh my gosh I get it now thank you both so much.

OpenStudy (anonymous):

Amplitude for any sinusoidal function, in words, can be described as the absolute value of the difference between the highest value the function goes through and lowest value, divided by 2.

OpenStudy (anonymous):

So the function y = 3sin(2x) + 5 has the highest value of 8 and the lowest value of 2, then the amplitude would be the absolute value of the difference of these highest and lowest values divided by two. It looks like this:\[\bf \frac{ \left[ 8-2 \right] }{ 2 }=3\]